4x^2 -100=0 I need this solved please.
Answer by stanbon(75887)

4x^2 -100=0
-----
Factor:
4(x^2-25) = 0
Divide by 4 to get:
x^2-25 = 0
----
Factor:
(x-5)(x+5) = 0
Solve:
x = 5 or x = -5
---------------------
Cheers,
Stan H.
-------------------